# Fragmentation in Collapsar Disks: Migration, Growth, and Emission

### v1

@@ Abstract
We present a parameter survey of fragmentation in collapsar disks, using a revised version of the Chen & Beloborodov (2007) model that determines the structure of steady state hyperaccretion disks in a general relativistic and neutrino cooled framework. We map out the range of disk conditions leading to gravitational instability alongside an exploration of the dimensionless cooling time $\beta$, which together determine whether fragmentation is likely to occur. We estimate the initial mass and density of fragments, finding that they occupy a unique region in the space of self-gravitating compact objects, with masses $M_{\rm f}\sim 10^{-3}M_{\odot}-10^{-1}M_{\odot}$ and densities $\rho_{\rm f}\sim 10^{8}-10^{11}\leavevmode\nobreak\ {\rm g\leavevmode\nobreak\ cm}^{-3}$. We then calculate their migration and mass growth (via Bondi-Hoyle accretion) as they inspiral through the collapsar disk. During a fragment’s migration to the central black hole, it can grow its mass up to a range $M_{\rm f}\sim 10^{-1}M_{\odot}-1M_{\odot}$. In most cases, the final fragment mass is larger than the minimum cold stable neutron star mass but much smaller than any observed neutron star. The fragment briefly achieves peak accretion rates comparable to (or even larger than) that of the central engine. We propose that these bound fragments may give rise to observable astrophysical phenomena, and we approximately model two of these: (i) gamma ray burst variability produced by a secondary, fragment-launched jet; (ii) the generation of non-vacuum gravitational waveforms accompanied by electromagnetic counterparts.
@@ 1 Introduction
Collapsar disks are extreme astrophysical environments, with accretion rates that can exceed $10^{14}$ times the naive (photon) Eddington limit (Di Matteo et al., 2002). The resulting densities in these disks lie in between the densities of white dwarfs and those of NSs, but the underlying microphysics differs from that of compact objects due to very high temperatures, which partially lift electron degeneracy and also power copious neutrino emission (Beloborodov, 2003). At the highest accretion rates ($\dot{M}\gtrsim 1M_{\odot}\leavevmode\nobreak\ {\rm s}^{-1})$, collapsar disks can develop the Toomre (Toomre, 1964) self-gravitational instability (Chen & Beloborodov, 2007).
@@
With these motivations, we perform a parameter survey of 1D, steady state general relativistic models for neutrino-cooled collapsar disks. Our work also explores subsequent interactions (migration and mass growth) between bound fragments and the disk they formed from. Our approach, which is based on the model of Chen & Beloborodov 2007, is much more approximate than a self-consistent magneto-hydrodynamic simulation, but is also far cheaper computationally, and allows us to explore the large parameter space of central masses, accretion rates, and disk microphysics. In §2, we introduce the underlying physics of the disk model (though details are elaborated in Appendices B-D). In §3, we present the primary results of our survey, focusing on which regions of parameter space exhibit sufficiently strong self-gravity and sufficiently short cooling times to form self-bound fragments. We also explore the evolution of these fragments due to orbital migration and mass growth through accretion. In §4, we make tentative estimates concerning the contribution of these fragments to GRB variability and GW production in collapsar disks, which are further discussed in §5. We conclude in §6.
@@ 2.3.1 Reduced Bondi-Hoyle accretion
Where $\sqrt{R_{{\rm BH}}R_{{\rm H}}}$ is the maximum impact parameter within a Hill sphere of radius $R_{{\rm H}}=r\left(\frac{m_{\rm f}}{3M}\right)^{1/3}$ (Murray & Dermott, 1999).
@@ 2.3.2 Fragment migration
The first torque we will consider is the “type I” torque arising from gravitational back-reaction: the gravity of the fragment linearly perturbs the collapsar disk, creating linear overdensities and underdensities, which in turn exert gravitational torques on the perturbing fragment (Goldreich & Tremaine, 1980). The torque exerted by the disk on the orbiting body usually causes it to lose energy and momentum, causing it to migrate inward. We write the type I torque as (Paardekooper et al., 2010):
@@
$\Gamma_{{\rm I/II}}$ =$=\begin{cases}\displaystyle\Gamma_{{\rm I}}&\kappa_{{\rm gap}}<20\\[8.61108pt] \displaystyle\Gamma_{{\rm II}}&\kappa_{{\rm gap}}\geq 20\end{cases},$ (19a)
$\dot{m}_{{\rm f}}$ =$=\begin{cases}\displaystyle\left(\frac{{\rm d}m_{{\rm f}}}{{\rm d}t}\right)_{{\rm RBH}}&\kappa_{{\rm gap}}<20\\[8.61108pt] \displaystyle\left(\frac{{\rm d}m_{{\rm f}}}{{\rm d}t}\right)_{{\rm RBHG}}&\kappa_{{\rm gap}}\geq 20.\end{cases}$ (19b)
Gas in accretion disks moves on quasi-circular orbits where attraction from the BH is balanced primarily by centrifugal force, but with a weak contribution from radial pressure gradients. The radial pressure gradient causes azimuthal gas velocities to be generically sub-Keplerian, creating another source of migratory torque. As the fragment accretes mass with sub-Keplerian specific angular momentum, it decreases the fragment’s specific angular momentum, leading to inward migration. The total torque exerted on the fragment from the gas [^4] Due to the uncertainties of fragment accretion and the collapsar environment, we neglect “thermal torques” arising from accretion feedback from the fragment on the collapsar disk (Masset, 2017; Grishin et al., 2024), though these can be important in other settings. is thus the sum of the type I/${\rm I/II}$ torque and the torque of the accumulated gas on the fragment:
@@ 3.1 Disk Stability
In this section we present our parameter survey across a wide range of central masses $M$, accretion rates $\dot{M}$ and effective viscosities $\alpha$. We have chosen physically motivated values for these parameters: for example, we consider the full range of black hole masses that are likely to result from stellar evolution, $3\lesssim M/M_{\odot}\lesssim 50$ (Spera & Mapelli, 2017), and a range of Shakura-Sunyaev viscosity coefficients, $0.30.01<\alpha<$ 0.3, that is broadly consistent with magnetohydrodynamic (MHD) accretion simulations (Penna et al., 2013; Siegel et al., 2019; Jiang et al., 2019). The accretion rates in collapsar disks remain uncertain, and multi-dimensional simulations of iron core collapse in rotating, massive stars find a large range of peak $\dot{M}$, spanning three orders of magnitude, from $0.01$ to $10M_{\odot}/{\rm s}$ (Ott et al., 2011; Sekiguchi & Shibata, 2011; Fujibayashi et al., 2022). As we shall see, these high accretion rates do not need to be sustained for the duration of the collapsar event but only for relatively short periods of time (long enough for a quasi-steady state disk solution to be established). We focus on higher accretion rates ($0.1M_{\odot}\leavevmode\nobreak\ {\rm s}^{-1}\leq\dot{M}\leq 10M_{\odot}\leavevmode\nobreak\ {\rm s}^{-1}$) because this is the portion of parameter space where gravitational instability can emerge (CB07). For the results shown in the main body of the paper we set the BHs spin to be $a=0.95$, however we have also recalculated some of our findings for a non-fiducidal, non-spinning case in Appendix F.4. As we will see, fragmentation almost always occurs at relatively large radii in collapsar disks where the effect of BH spin is relatively unimportant. The final evolution of fragments can be substantially more sensitive to spin, as this will affect how far inwards fragments can migrate in the disk, but we will discuss this issue later in §4.2.
@@ 3.1.2 Cooling time: $\beta$
For each run of our disk models, we calculated the radial profile of the dimensionless cooling time $\beta$ (Eq. 10) and looked for regions where both $Q<1$ and $\beta<35$ (for the rest of the main text, we set $\beta_{\rm c}=35$ as our definition of the critical fragmentation boundary, motivated by Chen et al. 2023; in Appendix F.3 we show more conservative results with $\beta_{\rm c}=10$). Fig. 3 shows $\beta$ profiles for four different disk solutions. In the selected examples, every disk model with $Q<1$ regions contains extensive regions of Toomre instability with $\beta<10$. It is less common for $\beta$ to go below the classic $\beta_{\rm c}=3$ value (Gammie, 2001), but in some of our models, particularly those with high accretion rates, it does (see e.g. the $M=3M_{\odot}$, $\dot{M}=10M_{\odot}\leavevmode\nobreak\ {\rm s}^{-1}$ example in Fig. 3).
Figure 3: Gammie’s $\beta$ parameter as a function of normalized radius, for the same four different disk solutions seen in Fig. 1. Lines are solid where $Q<1$ and dashed where $Q>1$. We shade two zones of $\beta$: in light-green, $3510<\beta<$ 35, and in light-cyan $\beta<10$, reflecting upper limits for fragmentation found in the simulations of Chen et al. (2023) and Brucy & Hennebelle (2021), respectively. The solutions that achieve Toomre instability (purple, green, red) possess $\beta<10$ sections of their $Q<1$ zones, although it is only in relatively extreme cases (e.g. the green $M=3M_{\odot}$, $\dot{M}=10M_{\odot}\leavevmode\nobreak\ {\rm s}^{-1}$) that $\beta$ falls below the classic $\beta_{\rm c}=3$ value (Gammie, 2001).
@@ 3.2 Fragment Initial Mass
Figure 5: Initial fragment mass $m_{\rm f}$ plotted against dimensionless radius $r/r_{\rm s}$ for a wide variety of different disk parameters $M$, ${\dot{m}}_{\rm*}=\dot{M}/(M_{\odot}\leavevmode\nobreak\ {\rm s}^{-1}$) and $\alpha$. Each string of points shows the regions of a given disk model where $Q\leq 1$ and the cooling parameter $\beta$ is underneath a critical value $\beta_{\rm c}$. Color coding indicates the value of $\beta_{\rm c}$ as well as the dominance (or not) of e- degeneracy pressure, as indicated in the legends. Grey dots indicate regions with $Q\leq 1$ but inefficient cooling ($\beta>35$) that likely prohibits fragmentation; black dots indicate regions where the disk models may fail due to heavy element production; magenta dots are placed at the innermost radii that can support fragmentation (we later choose these points to study fragment evolution). For large $M$, the instability regions are sometimes bifurcated; $Q>1$ intermediate zones are shown with dashed black lines. Top panel: results for fixed $\alpha=0.02$ and $\dot{M}=1M_{\odot}\leavevmode\nobreak\ {\rm s}^{-1}$. Middle panel: results for fixed $\alpha=0.02$ and $M=20M_{\odot}$. Bottom panel: results for fixed $M=20M_{\odot}$ and $\dot{M}=3M_{\odot}\leavevmode\nobreak\ {\rm s}^{-1}$. Fragment masses generally fall in the range $10^{-4}\lesssim m_{\rm f}/M_{\odot}\lesssim 10^{-1}$. While holding two of the other disk variables constant, the initial masses of fragments tend (depending on the appearance of bifurcated $Q<1$ regions) to increase with increasing $M$ or $\alpha$, or with decreasing $\dot{M}$.
@@
An important caution should be made concerning the physical conditions in the disk. At larger radii, the density and temperature ranges that follow from our model often fall in the typical range for outer envelopes (i.e. the crust) of proto-NSs and remnants of NS mergers (e.g. Haensel et al., 2007; Dehman et al., 2024). Typical equations of state of such matter predict that formation of heavy atomic nuclei with $Z\geqslant 3$ are energetically favorable. This option is ignored in the master Eqs. (1)—(3). However, to take heavy atomic nuclei into account, one has to consider a model of nuclear interactions, which is not a trivial task. Here we use the LS220 equation of state for hot dense matter (Lattimer & Swesty, 1991)888provided by the CompOSE database (Typel et al., 2015) https://compose.obspm.fr/. We use this model to explore qualitatively the possibility of heavy nuclei formation. Black dots in Fig. 5 mean that at a given $r$, the physical conditions in the disk ($\rho$, $T$, $Y_{e}$) are such that LS220 predicts more than a 5% mass fraction of heavy nuclei. This indicates that the present disk model may fail there. A comprehensive investigation of this problem is beyond the scope of the present work, but this post hoc check shows that in most cases, our neglect of heavy nuclei is only relevant at large radii well beyond $r_{\rm Q}$, and thus will not impact our estimates for fragment properties and evolution near the inner edge of the fragmentation zone.

### v2

@@ Abstract
We present a parameter survey of fragmentation in collapsar disks, using a revised version of the Chen & Beloborodov (2007) model that determines the structure of steady state hyperaccretion disks in a general relativistic and neutrino cooled framework. We map out the range of disk conditions leading to gravitational instability alongside an exploration of the dimensionless cooling time $\beta$, which together determine whether fragmentation is likely to occur. We estimate the initial mass and density of fragments, finding that they occupy a unique region in the space of self-gravitating compact objects, with masses $m_{\rm f,i}\sim 10^{-3}M_{\odot}-10^{-1}M_{\odot}$ and densities $\rho_{\rm f}\sim 10^{8}-10^{11}\penalty 10000\ {\rm g\penalty 10000\ cm}^{-3}$. We then calculate their migration and mass growth (via Bondi-Hoyle accretion) as they inspiral through the collapsar disk. During a fragment’s migration to the central black hole, it can grow its mass up to a range $m_{\rm f}\sim 10^{-1}M_{\odot}-1M_{\odot}$. In most cases, the final fragment mass is larger than the minimum cold stable neutron star mass but much smaller than any observed neutron star. The fragment briefly achieves peak accretion rates comparable to (or even larger than) that of the central engine. We propose that these bound fragments may give rise to observable astrophysical phenomena, and we approximately model two of these: (i) gamma ray burst variability produced by a secondary, fragment-launched jet; (ii) the generation of non-vacuum gravitational waveforms accompanied by electromagnetic counterparts.
@@ 1 Introduction
Collapsar disks are extreme astrophysical environments, with accretion rates that can exceed $10^{14}$ times the naive (photon) Eddington limit (Di Matteo et al., 2002). The resulting densities in these disks lie in between the densities of white dwarfs and those of NSs, but the underlying microphysics differs from that of compact objects due to very high temperatures, which partially lift electron degeneracy and also power copious neutrino emission (Beloborodov, 2003). At the highest accretion rates ($\dot{M}\gtrsim 1M_{\odot}\penalty 10000\ {\rm s}^{-1})$, collapsar disks can develop the Toomre (Toomre, 1964) self-gravitational instability (Chen & Beloborodov, 2007).
@@
With these motivations, we perform a parameter survey of 1D, steady state general relativistic models for neutrino-cooled collapsar disks. Our work also explores subsequent interactions (migration and mass growth) between bound fragments and the disk they formed from. Our approach, which is based on the model of Chen & Beloborodov 2007, is much more approximate than a self-consistent magneto-hydrodynamic simulation, but is also far cheaper computationally, and allows us to explore the large parameter space of central masses, accretion rates, and disk microphysics. In §2, we introduce the underlying physics of the disk model (though details are elaborated in Appendices A-D). In §3, we present the primary results of our survey, focusing on which regions of parameter space exhibit sufficiently strong self-gravity and sufficiently short cooling times to form self-bound fragments. We also explore the evolution of these fragments due to orbital migration and mass growth through accretion. In §4, we make tentative estimates concerning the contribution of these fragments to GRB variability and GW production in collapsar disks, which are further discussed in §5. We conclude in §6. Additional results that more fully explore the parameter space of our model are presented in the Supplementary Material (Appendix F).
@@ 2.2.2 Cooling time ($\beta$)
If the larger radii in our disks were to reach a gravitoturbulent state, they would likely self-regulate to the point of marginal stability with $Q\approx 1$. A similar type of self-regulation would occur if accretion feedback onto fragments heats the disk (Sirko & Goodman, 2003; Thompson et al., 2005); in this scenario, the disk could even be over-stabilized to a $Q>1$ state (Gilbaum & Stone, 2022). We do not consider the implications of self-regulation in this paper, deferring these questions for future work, but note that the impact of self-regulation will be minimized at the smallest radii in our Toomre-unstable zones.
@@ 2.3.1 Reduced Bondi-Hoyle accretion
Where $\sqrt{R_{{\rm BH}}R_{{\rm H}}}$ is the maximum impact parameter within a Hill sphere of radius $R_{{\rm H}}=r\left(\frac{m_{\rm f}}{3M}\right)^{1/3}$ (Murray & Dermott, 1999). These types of geometrical reductions in the naive Bondi-Hoyle cross-section have been tested and validated against numerical hydrodynamics simulations (Choksi et al., 2023; Li et al., 2023a).
@@ 2.3.2 Fragment migration
The first torque we will consider is the “type I” torque arising from gravitational back-reaction: the gravity of the fragment linearly perturbs the collapsar disk, creating linear overdensities and underdensities, which in turn exert gravitational torques on the perturbing fragment (Goldreich & Tremaine, 1980). The torque exerted by the disk on the orbiting body usually causes it to lose energy and momentum [^4] In this paper, we neglect the role of disk turbulence in generating stochasticity in the migration torques that we employ. Our treatment of migration is deterministic, but strong turbulence (from e.g. the gravitoturbulent $\beta\gg 1$ regime) can in some contexts lead to important stochasticity in migration (Johnson et al., 2006; Baruteau & Lin, 2010; Wu et al., 2024)., causing it to migrate inward. We write the type I torque as (Paardekooper et al., 2010):
@@
$\Gamma_{{\rm I/II}}$ $=\begin{cases}\displaystyle\Gamma_{{\rm I}}&\kappa_{{\rm gap}}<20\\[8.61108pt] \displaystyle\Gamma_{{\rm II}}&\kappa_{{\rm gap}}\geq 20\end{cases},$ (19a)
$\dot{m}_{{\rm f}}$ $=\begin{cases}\displaystyle\left(\frac{{\rm d}m_{{\rm f}}}{{\rm d}t}\right)_{{\rm RBH}}&\kappa_{{\rm gap}}<20\\[8.61108pt] \displaystyle\left(\frac{{\rm d}m_{{\rm f}}}{{\rm d}t}\right)_{{\rm RBHG}}&\kappa_{{\rm gap}}\geq 20.\end{cases}$ (19b)
We note that this reduced Bondi-Hoyle accretion rate is self-limited by the gap-opening process. Once the fragment exceeds a critical gap-opening mass, the depleted surface density in its annular region will greatly reduce its ability to grow further. Prior to gap opening, the reduced Bondi-Hoyle accretion rate can become quite large, in some cases even exceeding $\dot{M}$ by a factor of a few (Fig. 9), but fragment growth is quickly saturated by gap opening and $\dot{m}_{\rm f}$ can only exceed $\dot{M}$ for a few milliseconds (see Fig. 28).
Our estimate for the accretion rate assumes that annular gaps can only open due to nonlinear angular momentum exchange with surrounding gas via Lindblad resonances (i.e. the classic Type II picture, in which a massive perturber pumps angular momentum into nearby gas faster than it can be viscously refreshed). The unusual conditions in collapsar disks may motivate a second form of gap opening, one in which (reduced) Bondi-Hoyle accretion depletes surrounding gas faster than the viscous refresh time. While a full examination of these “depletion gaps” is beyond the scope of this work, we remark that it may be an important topic for future examination, as the emergence of such gaps would diminish final fragment masses.
Gas in accretion disks moves on quasi-circular orbits where attraction from the BH is balanced primarily by centrifugal force, but with a weak contribution from radial pressure gradients. The radial pressure gradient causes azimuthal gas velocities to be generically sub-Keplerian, creating another source of migratory torque. As the fragment accretes mass with sub-Keplerian specific angular momentum, it decreases the fragment’s specific angular momentum, leading to inward migration. The total torque exerted on the fragment from the gas [^5] Due to the uncertainties of fragment accretion and the collapsar environment, we neglect “thermal torques” arising from accretion feedback from the fragment on the collapsar disk (Masset, 2017; Grishin et al., 2024), though these can be important in other settings. is thus the sum of the type $\rm I/II$} torque and the torque of the accumulated gas on the fragment:
@@ 3.1 Disk Stability
In this section we present our parameter survey across a wide range of central masses $M$, accretion rates $\dot{M}$ and effective viscosities $\alpha$. We have chosen physically motivated values for these parameters: for example, we consider the full range of black hole masses that are likely to result from stellar evolution, $3\lesssim M/M_{\odot}\lesssim 50$ (Spera & Mapelli, 2017), and a range of Shakura-Sunyaev viscosity coefficients, $0.01<\alpha<0.3$, that is broadly consistent with magnetohydrodynamic (MHD) accretion simulations (Penna et al., 2013; Siegel et al., 2019; Jiang et al., 2019). The accretion rates in collapsar disks remain uncertain, and multi-dimensional simulations of iron core collapse in rotating, massive stars find a large range of peak $\dot{M}$, spanning three orders of magnitude, from $0.01$ to $10M_{\odot}/{\rm s}$ (Ott et al., 2011; Sekiguchi & Shibata, 2011; Fujibayashi et al., 2022). As we shall see, these high accretion rates do not need to be sustained for the duration of the collapsar event but only for relatively short periods of time (long enough for a quasi-steady state disk solution to be established). We focus on higher accretion rates ($0.1M_{\odot}\penalty 10000\ {\rm s}^{-1}\leq\dot{M}\leq 10M_{\odot}\penalty 10000\ {\rm s}^{-1}$) because this is the portion of parameter space where gravitational instability can emerge (CB07). For the results shown in the main body of the paper we set the BHs spin to be $a=0.95$, however we have also recalculated some of our findings for a non-fiducidal, non-spinning case in Appendix F.4. As we will see, fragmentation almost always occurs at relatively large radii in collapsar disks where the effect of BH spin is relatively unimportant. The final evolution of fragments can be substantially more sensitive to spin, as this will affect how far inwards fragments can migrate in the disk, but we will discuss this issue later in §4.2.
@@ 3.1.2 Cooling time: $\beta$
For each run of our disk models, we calculated the radial profile of the dimensionless cooling time $\beta$ (Eq. 10) and looked for regions where both $Q<1$ and $\beta<35$ (for the rest of the main text, we set $\beta_{\rm c}=35$ as our definition of the critical fragmentation boundary, motivated by Chen et al. 2023; in Appendix F.3 we show more conservative results with $\beta_{\rm c}=10$). Fig. 3 shows $\beta$ profiles for four different disk solutions. In the selected examples, every disk model with $Q<1$ regions contains extensive regions of Toomre instability with $\beta<10$. It is less common for $\beta$ to go below the classic $\beta_{\rm c}=3$ value (Gammie, 2001), but in some of our models, particularly those with high accretion rates, it does (see e.g. the $M=3M_{\odot}$, $\dot{M}=10M_{\odot}\penalty 10000\ {\rm s}^{-1}$ example in Fig. 3).
Figure 3: Gammie’s $\beta$ parameter as a function of normalized radius, for the same four different disk solutions seen in Fig. 1. Lines are solid where $Q<1$ and dashed where $Q>1$. We shade two zones of $\beta$: in light-green, $10<\beta<35$, and in light-cyan $\beta<10$, reflecting upper limits for fragmentation found in the simulations of Chen et al. (2023) and Brucy & Hennebelle (2021), respectively. The solutions that achieve Toomre instability (purple, green, red) possess $\beta<10$ sections of their $Q<1$ zones, although it is only in relatively extreme cases (e.g. the green $M=3M_{\odot}$, $\dot{M}=10M_{\odot}\penalty 10000\ {\rm s}^{-1}$) that $\beta$ falls below the classic $\beta_{\rm c}=3$ value (Gammie, 2001).
@@ 3.2 Fragment Initial Mass
Figure 5: Initial fragment mass $m_{\rm f}$ plotted against dimensionless radius $r/r_{\rm s}$ for a wide variety of different disk parameters $M$, ${\dot{m}}_{\rm*}=\dot{M}/(M_{\odot}\penalty 10000\ {\rm s}^{-1}$) and $\alpha$. Each string of points shows the regions of a given disk model where $Q\leq 1$ and the cooling parameter $\beta$ is underneath a critical value $\beta_{\rm c}$. Color coding indicates the value of $\beta_{\rm c}$ as well as the dominance (or not) of e- degeneracy pressure, as indicated in the legends. Grey dots indicate regions with $Q\leq 1$ but inefficient cooling ($\beta>35$) that likely prohibits fragmentation; black dots indicate regions where the disk models may fail due to heavy element production; magenta dots are placed at the innermost radii that can support fragmentation (we later choose these points to study fragment evolution). For large $M$, the instability regions are sometimes bifurcated; $Q>1$ intermediate zones are shown with dashed black lines. Top panel: results for fixed $\alpha=0.02$ and $\dot{M}=1M_{\odot}\penalty 10000\ {\rm s}^{-1}$. Middle panel: results for fixed $\alpha=0.02$ and $M=20M_{\odot}$. Bottom panel: results for fixed $M=20M_{\odot}$ and $\dot{M}=3M_{\odot}\penalty 10000\ {\rm s}^{-1}$. Fragment masses generally fall in the range $10^{-4}\lesssim m_{\rm f}/M_{\odot}\lesssim 10^{-1}$. While holding two of the other disk variables constant, the initial masses of fragments tend (depending on the appearance of bifurcated $Q<1$ regions) to increase with increasing $M$ or $\alpha$, or with decreasing $\dot{M}$.
@@
An important caution should be made concerning the physical conditions in the disk. At larger radii, the density and temperature ranges that follow from our model often fall in the typical range for outer envelopes (i.e. the crust) of proto-NSs and remnants of NS mergers (e.g. Haensel et al., 2007; Dehman et al., 2024). Typical equations of state of such matter predict that formation of heavy atomic nuclei with $Z\geqslant 3$ are energetically favorable. This option is ignored in the master Eqs. (1)—(3). However, to take heavy atomic nuclei into account, one has to consider a model of nuclear interactions, which is not a trivial task. Here we use the LS220 equation of state for hot dense matter (Lattimer & Swesty, 1991)999provided by the CompOSE database (Typel et al., 2015) https://compose.obspm.fr/. We use this model to explore qualitatively the possibility of heavy nuclei formation. Black dots in Fig. 5 mean that at a given $r$, the physical conditions in the disk ($\rho$, $T$, $Y_{e}$) are such that LS220 predicts more than a 5% mass fraction of heavy nuclei. This indicates that the present disk model may fail there. A comprehensive investigation of this problem is beyond the scope of the present work, but this post hoc check shows that in most cases, our neglect of heavy nuclei is only relevant at large radii well beyond $r_{\rm Q}$, and thus will not impact our estimates for fragment properties and evolution near the inner edge of the fragmentation zone.